Anton is twice as old as Marie. Eight years ago, Anton was eight years younger than three times Marie's age. How old is Marie now?
Answer by Maths68(1474)

Let
Present age of Anton = a
Present age of Marie = m

Anton is twice as old as Marie.
a=2m.........(1)
Eight years ago,
Anton age = a-8
Marie age = m-8
Then, Anton was eight years younger than three times Marie's age.
a-8=3(m-8)-8
a-8=3m-24-8
a=3m-24-8+8
a=3m-24.........(2)
Put the value of a from (1) to (2)
2m=3m-24
2m-3m=-24
-m=-24
Multiply by -1 both sides
m=24
Put the value of m in (1)
a=2m
a=2*24
a=48
Present age of Anton = a = 48 years old
Present age of Marie = m = 24 years old